Problem 7-17 Profitability Index versus NPV

Broxton Group, a consumer electronics conglomerate, is reviewing its annual budget in wireless technology. It is considering investments in three different technologies to develop wireless communication devices. Consider the following cash flows of the three independent projects. Assume the discount rate is 9 percent. Further, the company has only $29 million to invest in new projects this year.

Cash Flows (in $ millions)

Year L6 G5 Wi-Fi
0 −$ 8.0 −$ 21 −$ 29
1 12.0 19 27
2 8.5 34 41
3 5.5 29 29


a. Calculate the profitability index for each investment. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

Profitability index
L6
G5
Wi-Fi


b. Calculate the NPV for each investment. (Enter your answers in dollars, not millions of dollars. Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 1,234,567.89.)

NPV
L6 $
G5 $
Wi-Fi $
Answer #1

1.
=1+(-8+12/1.09+8.5/1.09^2+5.5/1.09^3)/8
=2.801307924

2.
=1+(-21+19/1.09+34/1.09^2+29/1.09^3)/21
=3.259125398

3.
=1+(-29+27/1.09+41/1.09^2+29/1.09^3)/29
=2.816304926

4.
=-8+12/1.09+8.5/1.09^2+5.5/1.09^3
=14410463.40

5.
=-21+19/1.09+34/1.09^2+29/1.09^3
=47441633.35

6.
=-29+27/1.09+41/1.09^2+29/1.09^3
=52672842.85

7.
Invest in L6 & G5